# Bypassing the Exponential Dependency: Looped Transformers Efficiently Learn In-context by Multi-step Gradient Descent

### v1

@@
Bo Chen bc7b@mtmail.mtsu.edu. Middle Tennessee State University.    Xiaoyu Li xli216@stevens.edu. Stevens Institute of Technology.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.
@@ 1 INTRODUCTION
Large Language Models (LLMs) have gained immense success and have been widely used in our daily lives, e.g., GPT4 [65], Claude 3.5 [5], and so on, based on its Transformer architecture [81]. One core emergent ability of LLMs is In-Context Learning (ICL) [14]. During ICL, the user provides an input sequence (prompts) containing some question-answer pairs as in-context examples and the goal query that the user cares about, where the examples and query are drawn from an unknown task. The LLMs can in-context learn from these examples and generate the correct answer for the goal query without any parameter update. Notably, these unknown tasks may never seen by LLMs during their pre-training and post-training, e.g., synthetic task [84]. Thus, many believe that the in-context learning mechanism is different from supervised learning or unsupervised learning, where the latter may focus on feature learning, while the ICL may perform some algorithm to learn. For instance, the Transformer can implement algorithm selection [12] or gradient descent [22, 80] by in-context learning forward pass.
@@
In Theorem 1.1, as long as the condition number is constant, we can see that the linear looped Transformer will perform better when the loop number is increasing, i.e., the error will exponentially decay to 00. Informally, for a small constant $\epsilon$, if we draw $X$ from Gaussian distribution, the condition number of $X^{\top}X$ will be $1\leq\kappa\leq 1+O(\epsilon)$ when $n\geq\Omega(d/\epsilon^{2})$. Thus, informally, we only need $O(d)$ numbers of in-context examples to guarantee a good performance. Furthermore, our preliminary experiments (Section 6) validate the above arguments and our theoretical analysis.
The main intuition of our analysis is that we find that a linear looped Transformer can explicitly perform gradient descent in its hidden states (Lemma 4.1 and Theorem 4.2). Thus, the error analysis can be directly solved by the standard convex optimization technique (Theorem 5.8).
Our contributions:
@@
Roadmap. This paper is structured as follows: we begin with a review of related work in Section 2, followed by essential definitions and foundational concepts in Section 3. Section 4 delves into the gradient computation analysis within the Looped Transformer architecture, examining both individual layer computations and the full looped structure. In Section 5, we analyze the error convergence of looped transformers under strong convexity and smoothness conditions, providing an upper bound on error after $T$ gradient descent iterations for linear vector generation. Section 6 presents our experimental results and findings. Finally, we conclude in Section 7.
@@ 2 RELATED WORK
Large Language Models.
@@
Researchers have developed numerous adaptation techniques to optimize LLM performance for specific applications. These include methods such as adapters [37, 101, 27, 72], calibration approaches [105, 102], multitask fine-tuning strategies [23, 95, 80, 96], prompt tuning techniques [24, 44], scratchpad approaches [64], instruction tuning methodologies [45, 17, 62], symbol tuning [85], black-box tuning [74], reinforcement learning from the human feedback (RLHF) [67], chain-of-thought reasoning [90, 43, 98, 103] and various other strategies. And here are more related works aiming at enhancing model efficiency without compromising performance, such as [18, 54, 49, 38, 86, 35, 92, 83, 55, 33, 36, 34, 51, 76, 73].
In-context Learning.
@@
A significant capability that has emerged from Large Language Models (LLMs) is in-context learning (ICL) [14, 88, 75]. This feature allows LLMs to generate predictions for new scenarios when provided with a concise set of input-output examples (referred to as a prompt) for a specific task, without requiring any modification to the model’s parameters. ICL has found widespread application across various domains, including reasoning [104], self-correction [69], machine translation [7] and many so on.
@@
Numerous studies have focused on enhancing the ICL and zero-shot capabilities of LLMs [63, 87, 82, 40]. A substantial body of research has been dedicated to investigating the underlying mechanisms of transformer learning [28, 93, 30, 41, 4, 59, 46, 8, 60, 78, 79, 99, 9, 94, 50, 47, 48, 56, 57] and in-context learning [22, 61, 71, 11, 2, 68, 58, 52, 6, 106, 100, 32, 15, 89, 91, 26, 70] through both empirical and theoretical approaches. Building upon these insights, our analysis extends further to elucidate ICL implementing multi-step gradient descent.
Looped Transformer.
@@
The concept of recursive inductive bias was first introduced by [21] into Transformers. Looping Transformers are also related to parameter-efficient weight-tying Transformers [16, 29] theoretically showed that the recursive structure of Looped Transformers enables them to function as Turing machines. [97] demonstrate that increasing the number of loop iterations can improve performance on some tasks. Recent studies [1, 31] have provided theoretical insights into the emulation capabilities of specific algorithms and their convergence during training, with a particular emphasis on in-context learning. However, the expressive capacity of Looped Transformers or Looped Neural Newrok [53, 42] in function approximation and their associated approximation rates remain largely unexplored territories.
@@ 3.1 Notations
For any integer $n$, we define $[n]=\{1,2,\dots,n\}$. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. For each $a,b\in\mathbb{R}^{n}$, we use $a\circ b\in\mathbb{R}^{n}$ to denote the Hardamard product, i.e. the $i$-th entry of $(a\circ b)$ is $a_{i}b_{i}$ for all $i\in[n]$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. For $n>d$, for any matrix $A\in\mathbb{R}^{n\times d}$, we use $\|A\|$ to denote the spectral norm of $A$, i.e. $\|A\|:=\sup_{x\in\mathbb{R}^{d}}\|Ax\|_{2}/\|x\|_{2}$. For a matrix $A$, we use $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ to denote the minimum and maximum eigenvalue of $A$, respectively.
@@ 3.2 In-context Learning
Assume $\theta^{*}$ uniformly draw form $d$ dimension unit sphere $\mathbb{S}^{d}$, where $\|\theta^{*}\|_{2}=1$. Let labels be $y:=X\theta^{*}\in\mathbb{R}^{n}$.
@@
Given in-context prompt examples/data $X\in\mathbb{R}^{n\times d}$ with their labels $y=X\theta^{*}$, where $\theta^{*}$ is unseen to model, and given a query $\alpha\neq 0\in\mathbb{R}$, the task is to output/generate a vector $q$ such that $\langle q,\theta^{*}\rangle$ is as close to $\alpha$ as possible. Thus, the prediction error is
$|\langle q,\theta^{*}\rangle-\alpha|.$
We remark that our vector generation task is as hard as the linear regression task, as they are dual problems. Solving any one of them requires to estimate the $\theta^{*}$.
Combining all above, we have the ICL prompt/input data for the model.
@@ 3.3 Linear Looped Transformer
where $M=\begin{bmatrix}\mathbf{0}_{n\times n}&\mathbf{0}_{n\times 1}\\ \mathbf{1}_{1\times n}&0\end{bmatrix}$ is a casual attention mask for text generation.
In Definition 3.5, we combine the query matrix and key matrix as $Q$ and combine the value matrix and output matrix as $P$ for simplicity following previous works [100, 29].
@@ 3.4 Linear Regression with Gradient Descent
In this section, we introduce some key concept of linear regression with gradient descent.
@@
Let $\widetilde{\theta}=(X^{\top}X)^{-1}X^{\top}y$ be the optimizer of $\ell(\theta)$.
@@
We define the condition number of input data as $\kappa:=\frac{\lambda_{\max}(X^{\top}X)}{\lambda_{\min}(X^{\top}X)}$.
@@ 4 GRADIENT COMPUTATION IN LOOPED TRANSFORMER
$=$ $~{}(M\circ\begin{bmatrix}XX^{\top}+yy^{\top}&Xq^{(0)}+\alpha y\\ q^{(0)\top}X^{\top}+\alpha y^{\top}&\theta^{(0)\top}q^{(0)}+\alpha^{2}\end{bmatrix})Z^{(0)}P$
@@
where the first step is from Definition 3.3 and Theorem 4.2, the second step follows $\|\theta^{*}\|_{2}=1$, the third step is from the linear properties of inner product, and the fourth step is from Cauchy-Schwarz inequality. ∎
@@
Let $\kappa$ be the condition number defined in Definition 3.11. Let $T$ be the number of loops. Let the initial point $q^{(0)}={\bf 0}_{d}$. Then, we have the final prediction error satisfies
@@
In their work, the linear looped transformer has an error bound ${8Td^{2}2^{2T}}/{\sqrt{n}}$, while our bound is $|\alpha|\cdot\exp(-\frac{T}{2\kappa})$. As the looped number $T$ increases, our error bounds will exponentially decay, while theirs is exponentially increase. Note that our linear vector generation task and their linear regression task are dual problems. Our results align with the common intuition that more steps of gradient descent lead to better performance.
@@ 5.1 Convexity and Smoothness Analysis
where the first step follows from Definition 3.8, and the second step follows from properties of norm. From Definition 5.2, we observe that
@@ 5.2 Main Result
Then substituting this back, we have

### v2

@@
Bo Chen bc7b@mtmail.mtsu.edu. Middle Tennessee State University.    Xiaoyu Li xiaoyu.li2@student.unsw.edu.au. University of New South Wales.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.
@@ 1 INTRODUCTION
Large Language Models (LLMs) have gained immense success and have been widely used in our daily lives, e.g., GPT4 [82], Claude 3.5 [5], and so on, based on its Transformer architecture [101]. One core emergent ability of LLMs is In-Context Learning (ICL) [14]. During ICL, the user provides an input sequence (prompts) containing some question-answer pairs as in-context examples and the goal query that the user cares about, where the examples and query are drawn from an unknown task. The LLMs can in-context learn from these examples and generate the correct answer for the goal query without any parameter update. Notably, these unknown tasks may never be seen by LLMs during their pre-training and post-training, e.g., synthetic task [104]. Thus, many believe that the in-context learning mechanism is different from supervised learning or unsupervised learning, where the latter may focus on feature learning, while the ICL may perform some algorithm to learn. For instance, the Transformer can implement algorithm selection [12] or gradient descent [31, 100] by in-context learning forward pass.
@@
In Theorem 1.1, as long as the condition number is constant, we can see that the linear looped Transformer will perform better when the loop number is increasing, i.e., the error will exponentially decay to 00. Informally, for a small constant $\epsilon$, if we draw $X$ from Gaussian distribution, the condition number of $X^{\top}X$ will be $1\leq\kappa\leq 1+O(\epsilon)$ when $n\geq\Omega(d/\epsilon^{2})$. Thus, informally, we only need $O(d)$ numbers of in-context examples to guarantee a good performance.
Furthermore, our preliminary experiments (Section 6) validate the above arguments and our theoretical analysis. The main intuition of our analysis is that we find that a linear looped Transformer can explicitly perform gradient descent in its hidden states (Lemma 4.1 and Theorem 4.2). Thus, the error analysis can be directly solved by the standard convex optimization technique (Theorem 5.8).
1.1 Our Contributions
@@
1.2 Roadmap
This paper is structured as follows: we begin with a review of related work in Section 2, followed by essential definitions and foundational concepts in Section 3. Section 4 delves into the gradient computation analysis within the Looped Transformer architecture, examining both individual layer computations and the full looped structure. In Section 5, we analyze the error convergence of looped transformers under strong convexity and smoothness conditions, providing an upper bound on error after $T$ gradient descent iterations for linear vector generation. Section 6 presents our experimental results and findings. Finally, we conclude in Section 7.
@@ 2 RELATED WORK
2.1 Large Language Models
@@
Researchers have developed numerous adaptation techniques to optimize LLM performance for specific applications. These include methods such as adapters [51, 123, 38, 90, 26, 15], calibration approaches [128, 124], multitask fine-tuning strategies [34, 116, 100, 117], prompt tuning techniques [35, 59], scratchpad approaches [81], instruction tuning methodologies [60, 20, 79], symbol tuning [105], black-box tuning [92], reinforcement learning from the human feedback (RLHF) [84], chain-of-thought reasoning [111, 58, 120, 126] and various other strategies.
And here are more related works aiming at enhancing model efficiency without compromising performance, such as [33, 87, 93, 39, 32, 66, 70, 72, 103, 106, 113, 52, 45, 46, 47, 48, 49, 50, 68, 64, 61, 96, 91, 94, 16, 22, 21, 23, 19, 24, 25, 27, 125, 57].
2.2 In-context Learning
@@
A significant capability that has emerged from Large Language Models (LLMs) is in-context learning (ICL) [14, 109, 95]. This feature allows LLMs to generate predictions for new scenarios when provided with a concise set of input-output examples (referred to as a prompt) for a specific task, without requiring any modification to the model’s parameters. ICL has found widespread application across various domains, including reasoning [127], self-correction [86], machine translation [7] and many others.
@@
Numerous studies have focused on enhancing the ICL and zero-shot capabilities of LLMs [80, 107, 102, 54]. A substantial body of research has been dedicated to investigating the underlying mechanisms of transformer learning [40, 114, 42, 55, 4, 76, 62, 8, 77, 98, 99, 121, 9, 115, 67, 63, 65, 73, 74] and in-context learning [31, 78, 89, 11, 2, 85, 75, 69, 6, 129, 122, 44, 17, 110, 112, 37, 88, 95, 108] through both empirical and theoretical approaches. Building upon these insights, our analysis extends further to elucidate ICL implementing multi-step gradient descent.
2.3 Looped Transformer
@@
The concept of recursive inductive bias was first introduced by [30] into Transformers. Looping Transformers are also related to parameter-efficient weight-tying Transformers [18, 41] theoretically showed that the recursive structure of Looped Transformers enables them to function as Turing machines. [119] demonstrates that increasing the number of loop iterations can improve performance on some tasks. Recent studies [1, 43] have provided theoretical insights into the emulation capabilities of specific algorithms and their convergence during training, with a particular emphasis on in-context learning.
However, the expressive capacity of Looped Transformers or Looped Neural Newrok [71, 56] in function approximation and their associated approximation rates remain largely unexplored territories. [118] studies enhancing autoregressive chain-of-thought through loop-aligned reasoning. In this work, we consider Looped Linear Transformers.
Table 1: There is a strict correspondence between the symbols we use and the common symbols of the general standard transformer attention. The specific choice of $Q$ and $P$ follows the setting in previous work [41, 44] for simplicity.
Meaning Dimension Standard Notation Our Notation
Input $n\times d$ $X$ $Z$
Query-key weight matrix $d\times d$ $W_{Q}W_{K}^{\top}$ $Q$
Value-output weight matrix $d\times d$ $W_{V}W_{O}^{\top}$ $P$
Query-key matrix $n\times n$ $QK^{\top}$ = $XW_{Q}W_{K}^{\top}X^{\top}$ $ZQX^{\top}Z$
@@ 3.1 Notations
For any integer $n$, we define $[n]=\{1,2,\dots,n\}$. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $e_{i}$ to denote a vector where only $i$-th coordinate is $1$, and other entries are 00. For each $a,b\in\mathbb{R}^{n}$, we use $a\circ b\in\mathbb{R}^{n}$ to denote the Hardamard product, i.e. the $i$-th entry of $(a\circ b)$ is $a_{i}b_{i}$ for all $i\in[n]$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $x_{i,j}$ to denote the $j$-th coordinate of $x_{i}\in\mathbb{R}^{n}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. For $n>d$, for any matrix $A\in\mathbb{R}^{n\times d}$, we use $\|A\|$ to denote the spectral norm of $A$, i.e. $\|A\|:=\sup_{x\in\mathbb{R}^{d}}\|Ax\|_{2}/\|x\|_{2}$. For a matrix $A$, we use $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ to denote the minimum and maximum eigenvalue of $A$, respectively.
@@ 3.2 In-context Learning
Assume $\theta^{*}$ uniformly draw form $d$ dimension unit sphere $\mathbb{S}^{d}$, where $\|\theta^{*}\|_{2}=1$. Let labels be
$y:=X\theta^{*}\in\mathbb{R}^{n}.$
@@
Given in-context prompt examples/data $X\in\mathbb{R}^{n\times d}$ with their labels $y=X\theta^{*}$, where $\theta^{*}$ is unseen to model, and given a query $\alpha\neq 0\in\mathbb{R}$, the task is to output/generate a vector $v$ such that $\langle v,\theta^{*}\rangle$ is as close to $\alpha$ as possible. Thus, the prediction error is
$|\langle v,\theta^{*}\rangle-\alpha|.$
We remark that our vector generation task is as hard as the linear regression task, as they are dual problems. Solving any one of them requires estimating the $\theta^{*}$.
Combining all of the above, we have the ICL prompt/input data for the model.
@@ 3.3 Linear Looped Transformer
where $M=\begin{bmatrix}\mathbf{0}_{n\times n}&\mathbf{0}_{n\times 1}\\ \mathbf{1}_{1\times n}&0\end{bmatrix}$ is a causal attention mask for text generation.
In Definition 3.5, we combine the query matrix and key matrix as $Q$ and combine the value matrix and output matrix as $P$ for simplicity following previous works [122, 41]. We refer readers to Table 1 for more details about connection between our notation and standard attention.
@@ 3.4 Linear Regression with Gradient Descent
In this section, we introduce some key concepts of linear regression with gradient descent.
@@
Let the optimizer of $\ell(\theta)$ be
$\widetilde{\theta}=(X^{\top}X)^{-1}X^{\top}y.$
@@
We define the condition number of input data as
$\kappa:=\frac{\lambda_{\max}(X^{\top}X)}{\lambda_{\min}(X^{\top}X)}.$
@@ 4 GRADIENT COMPUTATION IN LOOPED TRANSFORMER
$=$ $~{}(M\circ\begin{bmatrix}XX^{\top}+yy^{\top}&Xq^{(0)}+\alpha y\\ q^{(0)\top}X^{\top}+\alpha y^{\top}&q^{(0)\top}q^{(0)}+\alpha^{2}\end{bmatrix})Z^{(0)}P$
@@
$=$ $~{}\begin{bmatrix}\mathbf{0}_{n\times d}&\mathbf{0}_{n\times 1}\\ q^{(0)\top}X^{\top}X+\alpha y^{\top}X&q^{(0)\top}X^{\top}y+\alpha y^{\top}y\end{bmatrix}$
$~{}\cdot\begin{bmatrix}I_{d\times d}&\mathbf{0}_{d\times 1}\\ \mathbf{0}_{1\times d}&0\end{bmatrix}$
@@
Lemma 4.3 (Cauchy-Schwarz inequality).
Let $a,b\in\mathbb{R}^{d}$, we have
$|\langle a,b\rangle|\leq\|a\|_{2}\cdot\|b\|_{2}.$
@@
By Definition 3.3, we have $q=\mathsf{TF}(Z^{(0)};Q,P),\theta^{*}\rangle$. Then, we have
@@
where the first step is from Definition 3.3 and Theorem 4.2, the second step follows $\|\theta^{*}\|_{2}=1$, the third step is from the linear properties of inner product, and the fourth step is from Cauchy-Schwarz inequality (Lemma 4.3). ∎
@@
Let $X\in\mathbb{R}^{n\times d}$, $y\in\mathbb{R}^{n}$, and $\theta\in\mathbb{R}^{d}$ be defined in Definition 3.8. Let $\kappa$ be the condition number defined in Definition 3.11. Let $T$ be the number of loops. Let the initial point $q^{(0)}={\bf 0}_{d}$. Then, we have the final prediction error satisfies
@@
In their work, the linear looped transformer has an error bound ${8Td^{2}2^{2T}}/{\sqrt{n}}$, while our bound is $|\alpha|\cdot\exp(-\frac{T}{2\kappa})$. As the looped number $T$ increases, our error bounds will exponentially decay, while theirs increase exponentially. Note that our linear vector generation task and the linear regression task are dual problems. Our results align with the common intuition that more steps of gradient descent lead to better performance.
@@ 5.1 Convexity and Smoothness Analysis
where the first step follows from Definition 3.8, and the second step follows from properties of the norm. From Definition 5.2, we observe that
@@
Proof.
Let $\{\lambda_{i}\}_{i=1}^{d}$ be the eigenvalue of $X^{\top}X$, $\{v_{i}\}_{i=1}^{d}$ is the corresponding standard orthogonal feature vector, we have
$v^{\top}(X^{\top}X)v=$ $~{}(\sum_{i=1}^{d}c_{i}v_{i}^{\top})(X^{\top}X)(\sum_{j=1}^{d}c_{j}v_{j})$
$=$ $~{}\sum_{i=1}^{d}\sum_{j=1}^{d}c_{i}c_{j}v_{i}^{\top}(X^{\top}X)v_{j}$
$=$ $~{}\sum_{i=1}^{d}\sum_{j=1}^{d}c_{i}c_{j}\lambda_{j}v_{i}^{\top}v_{j}$
$=$ $~{}\sum_{i=1}^{d}c_{i}^{2}\lambda_{i}$
$\geq$ $~{}\lambda_{\min}(X^{\top}X)\sum_{i=1}^{d}c_{i}^{2}$
$=$ $~{}\lambda_{\min}(X^{\top}X)\|v\|_{2}^{2}$
where the first step follows from properties of symmetric matrix, the second step follows from the matrix being decomposed into linear combinations of eigenvectors, the third step follows from basic algebra, the fourth step follows from orthogonality of eigenvectors, the fifth step follows from $\lambda_{\min}$ denotes the minimum eigenvalue, the sixth step follows from the definition of norm. ∎
@@ 5.2 Main Result
Then, substituting this back, we have
@@
$=$ $~{}(1-1/\kappa)^{\kappa\cdot T/\kappa}\|\theta^{(0)}-\theta^{*}\|_{2}^{2}$